1) Suppose a​ five-year, $1,000 bond with annual coupons has a price of $ 902.95 and a yield to maturity of 6.1 %. What is the​ bond's coupon​ rate?

Can you also explain how to put this problem in the TVM solver for TI-83 plus
Answer #1

Using financial calculator
PV=-902.95
FV=1000
I/Y=6.1%
N=5
CPT PMT=37.90
Hence, coupon rate=37.90/1000=3.790%